The equation y = 5x + 5 estimates the number of U.S. households, in millions, expected to install devices that receive and manage broadband TV and Internet content to the home, where x is the number of years after 2002 (Source: Forrester Research). For whatyears will there be at least 20 million homes with these devices?
1
Expert's answer
2014-06-14T04:51:55-0400
For 2002, we have y=5*0+5=5 millions U.S. households. To find a year where will be at least 20 million homes with these devices we need to substitute y=20 million in the y=5x+5. We have 20=5x+5, or 5x=15, x=3. So, in 2002+3=2005 will be at least 20 million homes with these devices.

Answer. 2005